# Fast Inference Using Automatic Differentiation and Neural Transport in Astroparticle Physics

### v1

@@ Appendix C Ground Truth Bayesian Evidence and Contour for Synthetic Model
$\mathcal{Z}\equiv\int\ell(\boldsymbol{\theta};\boldsymbol{X})\mathop{{}\pi(\boldsymbol{\theta})}\mathop{}\!\mathrm{d}^{k}\boldsymbol{\theta}\,.$ (16)
@@
$\mathcal{Z}=\frac{1}{(2\pi)^{nk/2}}\frac{1}{(2\pi\tilde{\sigma}^{2})^{k}}\prod_{i=1}^{k}\iint e^{-n\mathcal{C}_{i}}e^{-\mathcal{E}_{i}(\mu_{i})}\mathop{}\!\mathrm{d}^{k/2}\boldsymbol{\mu}\mathop{}\!\mathrm{d}^{k/2}\boldsymbol{\mathcal{C}}\,,$ (17)
@@
$\int_{-\infty}^{\infty}e^{-\mathcal{E}_{i}(\mu_{i})}\mathop{}\!\mathrm{d}\mu_{i}\equiv\mathcal{I}_{i}\equiv\sqrt{\frac{\pi}{a_{i}}}\exp\left(\frac{b_{i}^{2}-4a_{i}c_{i}}{4a_{i}}\right)\,,$ (20)
@@
$\mathcal{Z}=\frac{1}{(2\pi)^{nk/2}}\frac{1}{(2\pi\tilde{\sigma}^{2})^{k}}\prod_{i=1}^{k}\int e^{-n\mathcal{C}_{i}}\mathcal{I}_{i}\mathop{}\!\mathrm{d}^{k/2}\boldsymbol{\mathcal{C}}\,.$ (21)
@@
$\mathcal{F}(\theta_{u},\theta_{v})\equiv\int\ell(\theta_{u},\theta_{v},\boldsymbol{\theta^{\prime}};\boldsymbol{X})\mathop{{}\pi(\theta_{u},\theta_{v},\boldsymbol{\theta^{\prime}})}\mathop{}\!\mathrm{d}^{k-2}\boldsymbol{\theta^{\prime}}\,,$ (22)
@@
$\frac{\int\mathcal{F}^{\prime}(\theta_{u},\theta_{v})\mathop{}\!\mathrm{d}\theta_{u}\mathop{}\!\mathrm{d}\theta_{v}}{\int\mathcal{F}(\theta_{u},\theta_{v})\mathop{}\!\mathrm{d}\theta_{u}\mathop{}\!\mathrm{d}\theta_{v}}=\rho\%\,,\quad\text{with}\quad\mathcal{F}^{\prime}(\theta_{u},\theta_{v})\equiv\begin{cases}\mathcal{F}(\theta_{u},\theta_{v})&\text{if}\quad\mathcal{F}(\theta_{u},\theta_{v})>\varsigma_{p\%}\\ 0&\text{if}\quad\mathcal{F}(\theta_{u},\theta_{v})\leq\varsigma_{p\%}\end{cases}\,.$ (23)
@@
Appendix D Effective Sample Size Calculation
@@
Appendix E Full Corner Plots
@@
Appendix F Bayesian Evidence from Nested Sampling Runs

### v2

@@ Appendix B Ground Truth Bayesian Evidence and Contour for Synthetic Model
$\mathcal{Z}\equiv\int\ell(\boldsymbol{\theta};\boldsymbol{X})\mathop{{}\pi(\boldsymbol{\theta})}\mathrm{d}^{k}\boldsymbol{\theta}\,.$ (16)
@@
$\mathcal{Z}=\frac{1}{(2\pi)^{nk/2}}\frac{1}{(2\pi\tilde{\sigma}^{2})^{k}}\prod_{i=1}^{k}\iint e^{-n\mathcal{C}_{i}}e^{-\mathcal{E}_{i}(\mu_{i})}\,\mathrm{d}^{k/2}\boldsymbol{\mu}\,\mathrm{d}^{k/2}\boldsymbol{\mathcal{C}}\,,$ (17)
@@
$\mathcal{I}_{i}\equiv\int_{-\infty}^{\infty}e^{-\mathcal{E}_{i}(\mu_{i})}\,\mathrm{d}\mu_{i}=\sqrt{\frac{\pi}{a_{i}}}\exp\left(\frac{b_{i}^{2}-4a_{i}c_{i}}{4a_{i}}\right)\,,$ (20)
@@
$\mathcal{Z}=\frac{1}{(2\pi)^{nk/2}}\frac{1}{(2\pi\tilde{\sigma}^{2})^{k}}\prod_{i=1}^{k}\int e^{-n\mathcal{C}_{i}}\mathcal{I}_{i}\,\mathrm{d}^{k/2}\boldsymbol{\mathcal{C}}\,.$ (21)
@@
$\mathcal{F}(\theta_{u},\theta_{v})\equiv\int\ell(\theta_{u},\theta_{v},\boldsymbol{\theta^{\prime}};\boldsymbol{X})\mathop{{}\pi(\theta_{u},\theta_{v},\boldsymbol{\theta^{\prime}})}\,\mathrm{d}^{k-2}\boldsymbol{\theta^{\prime}}\,,$ (22)
@@
$\frac{\int\mathcal{F}^{\prime}(\theta_{u},\theta_{v})\,\mathrm{d}\theta_{u}\,\mathrm{d}\theta_{v}}{\int\mathcal{F}(\theta_{u},\theta_{v})\,\mathrm{d}\theta_{u}\,\mathrm{d}\theta_{v}}=\rho\%\,,\quad\text{with}\quad\mathcal{F}^{\prime}(\theta_{u},\theta_{v})\equiv\begin{cases}\mathcal{F}(\theta_{u},\theta_{v})&\text{if}\quad\mathcal{F}(\theta_{u},\theta_{v})>\varsigma_{p\%}\\ 0&\text{if}\quad\mathcal{F}(\theta_{u},\theta_{v})\leq\varsigma_{p\%}\end{cases}\,.$ (23)
@@
Appendix C Effective Sample Size Calculation
@@
Appendix D Full Corner Plots
@@
Appendix E Bayesian Evidence from Nested Sampling Runs